# FANformer: Improving Large Language Models Through Effective Periodicity Modeling

### v1

@@ Abstract
Periodicity, as one of the most important basic characteristics, lays the foundation for facilitating structured knowledge acquisition and systematic cognitive processes within human learning paradigms. However, the potential flaws of periodicity modeling in Transformer affect the learning efficiency and establishment of underlying principles from data for large language models (LLMs) built upon it. In this paper, we demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. We introduce FANformer, which integrates Fourier Analysis Network (FAN) into attention mechanism to achieve efficient periodicity modeling, by modifying the feature projection process of attention mechanism. Extensive experimental results on language modeling show that FANformer consistently outperforms Transformer when scaling up model size and training tokens, underscoring its superior learning efficiency. To further validate the effectiveness of FANformer, we pretrain a FANformer-1B on 1 trillion tokens. FANformer-1B exhibits marked improvements on downstream tasks compared to open-source LLMs with similar model parameters or training tokens. The results position FANformer as an effective and promising architecture for advancing LLMs. Our code and pretrained model are available at https://github.com/YihongDong/FANformer.
[…]
[ tikzsetting= fill=blueish, skipabove=0.33em, skipbelow=0.33em, linewidth=1pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=2pt, innerbottommargin=2pt, linecolor=gray95, roundcorner=2pt, shadowsize=4pt, shadowcolor=gray95 ]answerbox
FANformer: Improving Large Language Models Through Effective Periodicity Modeling
Yihong Dong1, Ge Li1,2, Xue Jiang1, Yongding Tao1, Kechi Zhang1, Hao Zhu1, Huanyu Liu1, Jiazheng Ding2, Jia Li ♂1, Jinliang Deng3, and Hong Mei1,4 1School of Computer Science, Peking University 2aiXcoder 3The Hong Kong University of Science and Technology 4Advanced Institute of Big Data dongyh@stu.pku.edu.cn, lige@pku.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress across various natural language processing tasks, establishing themselves as a cornerstone of modern artificial intelligence Brown et al. (2020); Zhao et al. (2023); Minaee et al. (2024). The decoder-only Transformer architecture, in particular, has emerged as the de facto standard for LLM development due to its superior performance and scalability OpenAI (2023); DeepSeek-AI et al. (2024b); Groeneveld et al. (2024). Besides these advancements, Transformer-based models are also known for their immense demand for data and computational resources during training Kaplan et al. (2020); Hoffmann et al. (2022); Chowdhery et al. (2023). In comparison, humans are able to accomplish similar learning tasks with far fewer resources. This discrepancy suggests that existing LLM architectures still suffer from low learning efficiency, leaving substantial room for improvement in their ability to extract and generalize the knowledge from data.
@@
Figure 1: The performance of FANformer on periodicity modeling and language modeling. (a) shows the training loss of Transformer and FANformer on the fitting mod function with scalar input and their performance at the 4,000th epoch. (b) shows the average performance of FANformer-1B and the open-source LLMs with comparable model parameters and training tokens.
@@
Periodicity, characterized by recurring patterns, is a ubiquitous phenomenon in human life and learning processes Buzsaki (2006); Lake et al. (2017). The human brain leverages pattern recognition mechanisms to process information and acquire knowledge efficiently Zalta et al. (2020); Edalati et al. (2023); Zhan et al. (2018). However, general network architectures represented by Transformers have potential flaws in periodicity modeling, which could hinder their learning efficiency Dong et al. (2024b); Liu et al. (2020). As shown in Figure 1 (a), even for a simple mod function, Transformer demonstrates suboptimal performance despite being provided with sufficient training data and model capacity [^1] We sample 400K training data from the function of mod 5 and train a 110M parameter Transformer for 4K epochs.. This inefficiency can be exacerbated during the training process of LLMs to affect their performance, considering the periodicity hidden in large amounts of language data. Fourier Analysis Network (FAN) Dong et al. (2024b) has shown preliminary success in tasks with explicit or implicit periodic features, but its integration with Transformer architectures for large-scale language modeling remains an open challenge.
@@
In this paper, we present FANformer, a novel foundation architecture for LLMs that integrates FAN into the attention mechanism of Transformer to improve the learning efficiency and performance of LLMs, by achieving efficient periodicity modeling. It leverages FAN to introduce Fourier principles for capturing and representing periodic patterns, thereby enhancing the Transformer’s capability to learn and generalize from data. Specifically, we modify the feature projection process of attention mechanism to incorporate frequency-domain representations to facilitate capturing and modeling periodicity. Figure 1 (a) demonstrates the significant advantages of FANformer over Transformer on periodicity modeling, with faster convergence speed and better results. In Figure 1 (b), we can observe that FANformer-1B achieves superior performance with higher utilization efficiency of model parameter and training token when benchmarked against comparable Transformer-based LLMs.
@@
To comprehensively validate the effectiveness and scalability of FANformer, we conducted extensive experiments on language modeling tasks. The results of scaling both model parameters and training tokens highlight that FANformer consistently surpasses Transformer, requiring only 69.2% of model parameters or 79.7% of training tokens to achieve comparable performance. We also implement a complete pretraining pipeline to pretrain a 1.1-billion parameter FANformer (FANformer-1B) on 1 trillion tokens. Experiments on various downstream tasks demonstrate that FANformer-1B outperforms open-source LLMs of the same size with fewer training tokens, and exceeds LLMs with three times the parameters when using the same training token. Moreover, through further detailed analysis, we reveal that FANformer is a superior choice compared to other variant architectures and discover three interesting findings: 1) By observing the training process, we discover the notable enhancements in FANformer’s learning efficiency over Transformer as the model continues to learn from the data. 2) For larger FANformers, the optimal hyperparameter to control the proportion of periodicity modeling exhibits an increasing trend, suggesting that more powerful FANformers could extract richer hidden periodicity features. 3) FANformer facilitates the rule-based learning paradigm, effectively mitigating the occurrence of "holes" inherent in the case-based reasoning of Transformer Hu et al. (2024). These findings underscore the potential of FANformer as an effective and scalable architecture for advancing LLMs.
The main contributions of our work can be summarized as follows:
  • We first demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs.
@@
  • We propose FANformer, a novel LLM architecture, which uses a simple yet effective approach to integrate FAN into attention mechanism for efficient periodicity modeling, consistently outperforming Transformers in scaling model parameters and training tokens.
  • We pretrain and open-source FANformer-1B, which surpasses state-of-the-art publicly available LLMs with similar parameter counts or training token budgets on downstream tasks.
@@
2 Preliminary Knowledge
Fourier Analysis Network (FAN) Dong et al. (2024b) enhances neural networks by introducing Fourier principles for effective periodicity modeling. The core component of FAN is its layer design, which combines periodic basis functions with standard linear transformations. Given a input $\mathbf{X}$, the FAN layer is defined as:
$\text{FANLayer}(\mathbf{X})=$ $[\cos(W_{p}\mathbf{X})\|\sin(W_{p}\mathbf{X})$
$\quad\|\sigma(W_{\bar{p}}\mathbf{X}+B_{\bar{p}})]$ (1)
[…]
3 FANformer
In this section, we provide a detailed description of FANformer for sequence modeling and adopt a decoder-only model to illustrate the architecture.
Given an input sequence $\boldsymbol{s}=\left\{s_{1},s_{2},\cdots,s_{l}\right\}\in\mathbb{R}^{l}$, where $s_{i}$ denotes the $i$-th token and $l$ represents the length of sequence $\boldsymbol{s}$, it is first mapped to the input embedding as $\mathbf{X}^{0}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{l}]\in\mathbb{R}^{l\times d_{h}}$, where $d_{h}$ represents the model’s hidden dimension. The embedding is subsequently fed into the model to obtain the final output $\mathbf{X}^{N}$, with each $n$-th layer of FANformer processing $\mathbf{X}^{n-1}$, where $n\in[1,N]$. The core of each FANformer layer lies in a revised attention module that integrates FAN, referred to as the ATtention-Fourier (ATF) module.
@@ 3.1 ATF
$\mathbf{X}_{F}$ $=\text{FANLayer}^{\prime}(\mathbf{X}),$
$=[\cos(W_{p}x)||\sin(W_{p}x)||(B_{\bar{p}}+W_{\bar{p}}x)],$ (2)
where $\text{FANLayer}^{\prime}$ represents a variant of Eq. (1) with the activation function $\sigma$ in Eq. (1) replaced by the identity function, i.e., $\sigma(x)=x$, in this paper. On this basis, we employ the linear transform to $\mathbf{X}_{F}$ to compute QKV projections, i.e., $\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}\in\mathbb{R}^{l\times d_{h}}$, as follows:
@@
$\text{ATF}(\mathbf{X})=\text{softmax}\left(\frac{\mathbf{Q}_{F}\mathbf{K}_{F}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{F},$ (4)
@@
$\text{Head}_{i}=\text{ATF}(\mathbf{X}|\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i},\mathbf{W}_{V}^{i};k),$ (5)
@@
where $\mathbf{W}_{O}\in\mathbb{R}^{d_{h}\times d_{h}}$ is the learnable parameter of out projection matrix. Note that $\text{ATF}(\mathbf{X})$ is mathematically equivalent to $\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$ (the detailed derivations are provided in Appendix F). This equivalence enables a simple yet effective implementation of Multi-head ATF as shown in Figure 2, which can seamlessly incorporate various advancements in traditional attention mechanisms, such as FlashAttention Dao et al. (2022).
@@ 3.2 Overall Architecture
The FANformer model comprises $N$ stacked FANformer layers, where each FANformer layer consists of a Multi-head ATF module and a feed-forward network (FFN) module. Following the previous work Groeneveld et al. (2024); Touvron et al. (2023b), we employ SwiGLU Ramachandran et al. (2018); Shazeer (2020) and pre-norm Zhang et al. (2019) as the enhancements to Transformer-based LLMs. Specifically, the $n$-th FANformer layer can be defined as:
@@ 4 Evaluation
We begin with the implementation details of our experiments (Section 4.1), followed by a comprehensive evaluation of FANformer from three distinct perspectives: First, we investigate the scalability of FANformer by examining its performance trends on language modeling tasks with respect to model size and training tokens (Section 4.2). Second, we evaluate the capabilities of the pre-trained FANformer-1B model across multiple downstream tasks (Section 4.3). Third, we conduct an in-depth empirical analysis of FANformer, including ablation study, learning efficiency, hyperparameter impact, reasoning mechanism, etc. (Section 4.4). See Appendix A-E for more experiments.
@@ 4.1 Implementation Details
The experiments in this paper are conducted on eight GPUs of Tesla A100-PCIe-40G. We adopt the open language model OLMo Groeneveld et al. (2024) as the baseline Transformer. Building upon this foundation, we apply ATF into OLMo to construct our FANformer, where hyperparameter p is set to 0.25 by default. For pertaining FANformer-1B, we sample 1T training tokens from OLMo’s training data, i.e., Dolma Soldaini et al. (2024). For other experiments, we train models on a smaller sample of Dolma, i.e., Dolma v1_6-sample (AllenAI, 2023), with roughly 10B tokens. The detailed pretraining and experimental setups are provided in Appendix G.
Table 1: Zero-shot performance of FANformer-1B versus other comparable open-source LLMs on 8 core tasks from the downstream evaluation suite following OLMo Groeneveld et al. (2024). The results of baselines are taken from the work Groeneveld et al. (2024); Ye et al. (2024); Dong et al. (2024a).

### v2

@@ Abstract
Periodicity, as one of the most important basic characteristics, lays the foundation for facilitating structured knowledge acquisition and systematic cognitive processes within human learning paradigms. However, the potential flaws of periodicity modeling in Transformer affect the learning efficiency and establishment of underlying principles from data for large language models (LLMs) built upon it. In this paper, we demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. We introduce FANformer, which adapts Fourier Analysis Network (FAN) into attention mechanism to achieve efficient periodicity modeling, by modifying the feature projection process of attention mechanism. Extensive experimental results on language modeling show that FANformer consistently outperforms Transformer when scaling up model size and training tokens, underscoring its superior learning efficiency. Our pretrained FANformer-1B exhibits marked improvements on downstream tasks compared to open-source LLMs with similar model parameters or training tokens. Moreover, we reveal that FANformer exhibits superior ability to learn and apply rules for reasoning compared to Transformer. The results position FANformer as an effective and promising architecture for advancing LLMs. Our code and pretrained model are available at https://github.com/YihongDong/FANformer.
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress across various natural language processing tasks, establishing themselves as a cornerstone of modern artificial intelligence (Brown et al., 2020; Zhao et al., 2023; Minaee et al., 2024). The decoder-only Transformer architecture, in particular, has emerged as the de facto standard for LLM development due to its superior performance and scalability (OpenAI, 2023; DeepSeek-AI et al., 2024a; Groeneveld et al., 2024). Besides these advancements, Transformer-based models are also known for their immense demand for data and computational resources during training (Kaplan et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023). In comparison, humans are able to accomplish similar learning tasks with far fewer resources. This discrepancy suggests that existing LLM architectures still suffer from low learning efficiency, leaving substantial room for improvement in their ability to extract and generalize the knowledge from data.
@@
Figure 1: The performance of FANformer on periodicity modeling and language modeling. (a) shows the training loss of Transformer and FANformer on the fitting mod function and their performance at the 4,000th epoch. (b) shows the average performance of 8 core commonsense tasks between FANformer-1B and the open-source LLMs with comparable model parameters and training tokens.
@@
Periodicity, characterized by recurring patterns, is a ubiquitous phenomenon in human life and learning processes (Buzsaki, 2006; Lake et al., 2017). The human brain leverages pattern recognition mechanisms to process information and acquire knowledge efficiently (Zalta et al., 2020; Edalati et al., 2023; Zhan et al., 2018). However, general network architectures represented by Transformers have potential flaws in periodicity modeling, which could hinder their learning efficiency (Dong et al., 2024a; Liu et al., 2020). As shown in Figure 1 (a), even for a simple mod function, Transformer demonstrates suboptimal performance despite being provided with sufficient training data and model capacity [^1] We sample 400K training data from the function of mod 5 and train a 110M Transformer for 4K epochs.. This inefficiency can be exacerbated during the training process of LLMs to affect their performance, considering the periodicity hidden in large amounts of language data. Fourier Analysis Network (FAN) (Dong et al., 2024a) has shown preliminary success in tasks with explicit or implicit periodic features, but its adaptation to Transformer architectures for large-scale language modeling remains an open challenge.
@@
In this paper, we present FANformer, a novel foundation architecture for LLMs that adapts FAN into the attention mechanism of Transformer to improve learning efficiency and model performance through effective periodicity modeling. It leverages FAN to introduce Fourier principles for capturing and representing periodic patterns, thereby enhancing the Transformer’s capability to learn and generalize from data. Specifically, we modify the feature projection process of attention mechanism to incorporate frequency-domain representations to facilitate modeling periodicity. Figure 1 (a) demonstrates the significant advantages of FANformer over Transformer on periodicity modeling, with faster convergence speed and better results. In Figure 1 (b), we can observe that FANformer-1B achieves superior performance with higher utilization efficiency of model parameters and training tokens when benchmarked against comparable Transformer-based LLMs.
@@
To comprehensively validate the effectiveness and scalability of FANformer, we conduct extensive experiments on language modeling tasks. The results of scaling both model parameters and training tokens highlight that FANformer consistently surpasses Transformer, requiring only 69.2% of model parameters or 79.7% of training tokens to achieve comparable performance. We also implement a complete pretraining pipeline to pretrain a 1.1-billion parameter FANformer (FANformer-1B) on 1 trillion tokens. Experiments on various downstream tasks demonstrate that FANformer-1B outperforms open-source LLMs of the same size with fewer training tokens, and exceeds LLMs with three times the parameters when using the same training token. Through further analysis, we reveal that FANformer is a superior choice compared to other variant architectures and discover three interesting findings: 1) By observing the training process, we discover the notable enhancements in FANformer’s learning efficiency over Transformer as the model continues to learn from the data. 2) FANformer facilitates the rule-based reasoning paradigm, mitigating the occurrence of "holes" inherent in the case-based learning of Transformer (Hu et al., 2024). Under the stress test of logical reasoning (Wang et al., 2024), FANformer-1B demonstrates superior performance compared to OLMo-1B and Qwen2.5-1.5B. 3) FANformer’s representational capacity consistently surpasses that of Transformer across various layer depths, as evidenced by evaluations of the model’s Lipschitz constant (Latorre et al., 2020). These findings underscore the potential of FANformer as an effective and scalable architecture for advancing LLMs.
@@
The main contributions of our work can be summarized as three points: ❶ We first demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. ❷ We propose FANformer, a novel LLM architecture, which uses a simple yet effective approach to adapt FAN into attention mechanism for efficient periodicity modeling, consistently outperforming Transformers in scaling model parameters and training tokens. ❸ We pretrain and open-source FANformer-1B, which surpasses SOTA publicly available LLMs with similar parameter counts or training token budgets on downstream tasks.
2 Motivation
In this section, we combine formalization with illustrative cases to demonstrate why periodicity modeling facilitates language modeling and reasoning, thereby elucidating the motivation behind the development of FANformer.
The essence of periodicity lies in the repetitive manifestation of certain invariance under transformations, which can be strictly defined through invariance under group actions in Abstract Algebra Dummit and Foote (2004). Let $X$ be a set and $G$ be a group acting on $X$. An element $x\in X$ is said to be periodic with respect to the action of $G$ if there exists a non-identity element $p\in G$ such that $p\cdot x=x$, where $\cdot$ denotes the group action. The element $p$ is called a period of $x$. Periodicity implies that $x$ is invariant under the action of the cyclic subgroup generated by $p$, denoted by $\langle p\rangle$. For example, $f(a)=f(a+T)$ can be seen as a specific instance of the abstract definition $p\cdot x=x$, where $x=f$, $p=T$, and the group action is translation. When the input $a$ and the group $G$ are extended to higher dimensions or non-temporal domains, the manifestation of the period $T$ also changes accordingly. Crucially, for many reasoning tasks, the underlying operation or inference rule remains invariant across structurally similar subproblems, that is, for all inputs belonging to a certain equivalence class, the same functional rule is applied, which precisely reflects periodic invariance.
Consider addition as an illustrative case: let $f$ represent the addition operation, $a$ denote the digit position index, and let the period $T=1$ correspond to the positional shift in place value. The reasoning proceeds as:
Example: 357 + 286 = ?
Digit-wise operations: units (7 + 6 = 13 $\rightarrow$ write 3, carry 1); tens (5 + 8 + 1 = 14 $\rightarrow$ write 4, carry 1); hundreds (3 + 2 + 1 = 6 $\rightarrow$ write 6).
Result: 643
Thus, the periodicity of addition is manifested in the repeated application of the same rule across different digit positions, where the rule itself remains invariant under positional shifts. It can be extended to other reasoning, such as logical reasoning, and the scenario is analogous: when a neural network extracts a feature applicable to specific conditions or premises, it repeatedly applies the same invariant rules across analogous subproblems. Such periodic invariance allows for enhancing both the learning efficiency and generalization capability of neural models by reducing redundancy and reinforcing conceptual regularities.
3 FANformer
We will provide a detailed description of FANformer for sequence modeling and adopt a decoder-only model to illustrate the architecture.
Given an input sequence $\boldsymbol{s}=\left\{s_{1},s_{2},\cdots,s_{l}\right\}$, where $s_{i}$ denotes the $i$-th token and $l$ represents the length of sequence $\boldsymbol{s}$, it is first mapped to the input embedding as $\mathbf{X}^{0}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{l}]\in\mathbb{R}^{l\times d_{h}}$, where $d_{h}$ represents the model’s hidden dimension. The embedding is subsequently fed into the model to obtain the final output $\mathbf{X}^{N}$, with each $n$-th layer of FANformer processing $\mathbf{X}^{n-1}$, where $n\in[1,N]$. The core of each FANformer layer lies in a revised attention module that incorporates a modified FAN layer, referred to as the ATtention-Fourier (ATF) module.
@@ 3.1 ATF
$\mathbf{X}_{F}=\text{FANLayer}^{\prime}(\mathbf{X})=[\cos(W_{p}\mathbf{X})||\sin(W_{p}\mathbf{X})||(W_{\bar{p}}\mathbf{X}+B_{\bar{p}})],$ (1)
where $\text{FANLayer}^{\prime}$ represents a variant of the original FAN layer (i.e., Eq. (10)) with the activation function $\sigma$ in Eq. (10) replaced by the identity function, i.e., $\sigma(x)=x$, in this paper, and hyperparameter $p$ is defined as the proportion of $\frac{d_{W_{p}}}{d_{h}}$. On this basis, we employ the linear transform to $\mathbf{X}_{F}$ to compute QKV projections, i.e., $\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}\in\mathbb{R}^{l\times d_{h}}$, as follows:
@@
$\text{ATF}(\mathbf{X}|\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V})=\text{softmax}\left(\frac{\mathbf{Q}_{F}\mathbf{K}_{F}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{F},$ (3)
@@
$\text{Head}_{i}=\text{ATF}(\mathbf{X}|\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i},\mathbf{W}_{V}^{i}),$ (4)
@@
where $\mathbf{W}_{O}\in\mathbb{R}^{d_{h}\times d_{h}}$ is the learnable parameter of out projection matrix. Note that $\text{ATF}(\mathbf{X})$ is mathematically equivalent to $\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$ (the detailed derivations are provided in Appendix I). This equivalence enables a simple yet effective implementation of Multi-head ATF as shown in Figure 2, which can seamlessly incorporate various advancements in traditional attention mechanisms, such as FlashAttention (Dao et al., 2022).
@@ 3.2 Overall Architecture
The FANformer model comprises $N$ stacked FANformer layers, where each FANformer layer consists of a Multi-head ATF module and a feed-forward network (FFN) module. Following the previous work (Groeneveld et al., 2024; Touvron et al., 2023a), we employ SwiGLU (Ramachandran et al., 2018; Shazeer, 2020) and pre-norm (Zhang and Sennrich, 2019) as the enhancements to Transformer-based LLMs. Specifically, the $n$-th FANformer layer can be defined as:
@@ 4 Evaluation
We begin with the implementation details of our experiments (Section 4.1), followed by a comprehensive evaluation of FANformer from three distinct perspectives: First, we investigate the scalability of FANformer by examining its performance trends on language modeling tasks with respect to model size and training tokens (Section 4.2). Second, we evaluate the capabilities of the pre-trained FANformer-1B model across multiple downstream tasks (Section 4.3). Third, we conduct an in-depth empirical analysis of FANformer, including ablation study, learning efficiency, reasoning mechanism, representational capacity, and more (Section 4.4). See Appendix A-G for more experiments.
@@ 4.1 Implementation Details
The experiments are conducted on 80 A100 GPUs. We build FANformer upon the foundation of OLMo (Groeneveld et al., 2024), as it provides a solid pretraining framework of LLMs, with the hyperparameter $p$ set to 0.25 by default. For pretraining FANformer-1B, we randomly sample 1T training tokens from OLMo’s training data, i.e., Dolma (Soldaini et al., 2024). For other experiments, we train models on a smaller sample of Dolma, i.e., Dolma v1_6-sample (AllenAI, 2023), with roughly 10B tokens. The detailed pretraining and experimental setups are provided in Appendix J.
4.2 Scalability of FANformer
Setup.
We follow OLMo’s configuration and vary the FFN’s intermediate dimension $d_{f}$ to keep the number of parameters consistent for all models in this experiment. For scaling up model parameters, we adopt Dolma v1_6-sample as training data and train LLMs from 268M to 7B. We compare FANformer with the standard Transformer and a variant of FANformer, termed Transformer+ATM, which uses MLP layer instead of FAN layer in FANformer. For scaling up training tokens, we train 1B LLMs on the first 200 billion of our sampled 1T tokens.
Figure 3: Language modeling loss of scaling up model parameters and training tokens. Left: we train LLMs from 268M to 7B parameters. Right: we evaluate LLMs every 20B tokens up to 200B tokens.
Results.
As shown in Figure 3, the scaling law (Kaplan et al., 2020) empirically aligns well with the results obtained from our FANformer, underscoring its superior scalability properties. Figure 3 (left) reveals that the implementation of FAN consistently surpasses the performance of the standard Transformer across a range of model sizes. This finding highlights FANformer’s enhanced scalability in terms of parameter efficiency, as it achieves comparable performance with only 69.2% of the parameters required by the standard Transformer. Notably, the scaling curve of Transformer+ATM closely overlaps with that of the standard Transformer, indicating that merely revising attention mechanisms using MLP Layer is insufficient. This observation further emphasizes that FANformer’s performance gains are not attributable to network depth increase, but rather to its special architectural design. Figure 3 (right) demonstrates that FANformer achieves performance parity with the standard Transformer while utilizing significantly fewer training tokens. Specifically, FANformer requires only 159.6B training tokens to match the performance of the standard Transformer trained on 200B tokens, representing a 20.3% reduction in training resource requirements. These findings suggest that FANformer exhibits superior utilization efficiency in terms of both model parameters and training tokens compared to the standard Transformer architecture.
4.3 Performance of FANformer-1B
We pretrain FANformer-1B on 1 trillion tokens and report zero-shot performance on a set of commonsense downstream tasks, following previous work (Brown et al., 2020; Touvron et al., 2023b; Groeneveld et al., 2024, inter alia).
Setup.
The downstream evaluation suite consists of 8 core commonsense tasks, including ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SCIQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2020). We compare pretrained FANformer-1B to seven open-source LLMs with comparable model parameters or training tokens, including Qwen2.5-1.5B (Team, 2024), R1-Distill-Qwen1.5B (DeepSeek-AI et al., 2025), Llama-3.2-1B (Dubey et al., 2024), TinyLlama-v1.1-1B (Zhang et al., 2024), OLMo-1B (Groeneveld et al., 2024), OpenLLaMA-v2-3B (Geng and Liu, 2023), and StableLM-v2-3B (Tow, 2023).
Table 1: Zero-shot performance of FANformer-1B versus other comparable open-source LLMs on 8 core tasks from the downstream evaluation suite following OLMo. The results of baselines are taken from the previous works (Groeneveld et al., 2024; Ye et al., 2024; Dong et al., 2024b).